you are doing a statistical experiment where you roll a die and pick a color. the color choices are blue, green, or purple. Give the sample space for this experiment. What is the probability that the outcome of the experiment is:
a) a purple and a 3?
b) a blue and an even number?
c) a green or a 5?
I asked this before but didn't get the answers for sample space just probabilty. I made that mistake and the teacher told me that it was wrong.
I greatly appreciate all your help with this. thank you!

Answer by ewatrrr(24785)


Hi,
Giving the Sample space as follows:
BLUE |_1__2__3__4__5__6
GREEN |_1__2__3__4__5__6
PURPLE |_1__2__3__4__5__6
Sample Space with 18 Possible Outcomes:
P(a purple and a 3) = 1/18
P( a blue and an even number) = 3/18
P( a green or a 5) = 6/18 + 3/18 - 1/18 = 8/18